# Aggregation Model Hyperparameters Matter in Digital Pathology

### v2

@@ 3.4 Evaluation Datasets
We assess the performance of feature extraction models in binary classification across three distinct digital pathology datasets. This approach guarantees that our findings are robust and not confined to a single dataset, thereby providing a more generalizable answer to our research questions. These datasets consist of WSIs of H&E-stained histopathology slides. On each dataset we evaluate the performance of every feature extractor using all the possible hyperparameter sets for the aggregation model being used. This setup results in 81 different hyperparameter sets for 7 feature extraction models, 2 aggregation models and 3 datasets, leading to a comparison of $81\times 7\times 2\times 3=3402$ experiments.
In evaluating the binary classification performance of each configuration, we utilized the area under the receiver operating characteristic curve (AUROC) as well as average precision (AP). Both AUROC and AP range from 0 to 1. A higher AUROC score indicates that the model has a better capability to distinguish between the positive and negative classes, thus correctly identifying true positives and true negatives. A higher AP score shows that the model more accurately predicts positive instances across all levels of recall, effectively balancing precision and recall. The metric for every evaluation on the test set is obtained by using the aggregation model parameters at the epoch with highest validation score during training.
COO: Binary classification of cell of origin (COO). Each image contains the COO prediction label of activated B-cell like (ABC) or germinal center B-cell like (GCB) tumors in diffuse large B-cell lymphoma (DLBCL). 709 WSIs from two internal datasets were used. This data closely mirrors real-world data, since it is crucial to assess classification approaches in digital pathology using such data and tasks. The WSIs (40$\times$ magnification) have been scanned by Ventana DP200 scanners. The artifact-free tissue tiles of this dataset were combined and randomly split into 70% training set, 15% validation set and 15% test set.
@@
Figure 3: The heatmap shows the performance of every aggregation model hyperparameter set for each feature extraction model. The lighter the colour the higher the AUROC. The heatmaps are shown for all three datasets and for both AttentionMIL and TransMIL approaches. The red colored legend shows how the hyperparameter sets are ordered on the heatmap. For example, the lowest learning rate is used for the first third of the hyperparameter sets, followed by the next higher learning rate. The legend should facilitate detecting hyperparameters sets with consistently high performance. However, no clear patterns are visible in the heatmaps indicating that aggregation model hyperparameters need to be accounted for when comparing feature extraction models.
Figure 4: Here, we compare seven feature extraction models while accounting for different aggregation model hyperparameter sets. Specifically, each boxplot represents the performance range when fixing the feature extraction model by using 81 different hyperparameter sets for the aggregation model. We show this for three different datasets and for both AttentionMIL and TransMIL approaches. From the boxplots it can be seen that the performance fluctuation is large as well as the overlap between different feature extraction models.
@@ 4 Results
In this section, we answer the questions posed in Section 3. First, we examine if an optimal aggregation model hyperparameter set exists, which is optimal for several feature extraction models. Subsequently, we compare the performance of different feature extraction models while considering the fluctuation of performance observed for different aggregation model hyperparameters.
4.1 Aggregation Model Influence
Hyperparameters play a crucial role in machine learning. Thus, it is expected to encounter classification performance fluctuations when changing the aggregation model hyperparameters. We herein examine how large the performance fluctuation is and if there is a co-dependency between the feature extraction models and aggregation model hyperparameters. The latter is already qualitatively confirmed by the motivating Figure 1, where we observe that the ranking for the best feature extractor model changes based on the chosen aggregation model hyperparameter set.
Figure 3 offers more insight into the effect of different hyperparameters by showing the classification performance for every hyperparameter set and feature extraction model as a heatmap. This heatmap, showing the AUROC score, shows similar trends as the one with the AP score shown in the supplementary. The legend above the heatmaps is included to facilitate detecting patterns in the heatmaps. For example the hyperparameter sets with the lowest learning rate are to the left of the heatmap, which are seen to produce worse performance in the case of the COO dataset with CTransPath. From this figure we can conclude two things.
No optimal hyperparameter: There does not seem to be a set of hyperparameters for the aggregation model that leads to superior results across feature extraction models. This would be clearly visible with a bright column spanning across all the feature extractor models.
Dataset-specific hyperparameters: Even when examining the best aggregation model hyperparameter set for a single feature extraction model no clear trend is visible. For some datasets and feature extraction models, the optimal learning rate seems to be important when using AttentionMIL. For TransMIL the number of attention blocks are more relevant. This can be seen by the patterns of the legend being replicated in the heatmaps. However, for most datasets and feature extraction models, no clear pattern from the color-coded hyperparameters are replicated in the heatmaps. This indicates that even for a single feature extractor model the best performing MIL hyperparameter set is not obvious and requires investigation.
Given the limitation of finding an optimal hyperparameter set, a distribution of hyperparameters should be considered when comparing feature extraction models. Otherwise, the improved performance of one feature extraction model over another could solely be due to the chosen aggregation model hyperparameter set.
@@ 4.2 Feature Extractor Comparison
As seen above, when developing new feature extractor models, it is crucial to carefully evaluate the performance gain compared to previous models. In Figure 4, we visualize the performance of feature extractor models across different datasets. Instead of showing a single result for a fixed aggregation model hyperparameter set, which is common practice in literature, we show the results of the feature extractor model over all 81 hyperparameter sets. We use box plots to effectively visualize the corresponding metrics. The performance variation linked to different aggregation model hyperparameters, previously observed in Figure 3, is reflected here in the vertical spread of the box plots. The large range of the boxes indicate that the overlap of the performance between different feature extractors is large. Remarkably, there are some clear trends to be noted despite these large ranges of results.
Training on DP datasets is necessary: The DINOv2 model trained on natural images performs poorly compared to all other models, consistently for all the datasets. From this we can conclude that it is indeed important to train feature extraction models on digital pathology datasets.
@@
5 Conclusion
In this study, we challenge the prevailing methodology for comparing feature extractor models in digital pathology literature, demonstrating that it may yield misleading results. Our investigation highlights that relying solely on a single aggregation model hyperparameter set can disproportionately favor certain feature extractor models while disadvantaging others. Our comprehensive analysis, taking into account performance variations across multiple hyperparameter sets of the aggregation model, reveals a considerable overlap in performance between different feature extractor models. Significantly, we find no universal aggregation model hyperparameter set that is uniformly effective for all feature extractors. Our work is limited though by only looking at classification tasks. In addition, the DINOv2 model we trained on digital pathology images might be subpar to other models due to computational and dataset limitations. Nevertheless, we believe this work will contribute to a more nuanced evaluation of feature extractor models that will help gain insight and further accelerate this rapidly evolving field.
Supplementary Material
Figure 5: Here, we use the average precision (AP) metric to compare seven feature extraction models while accounting for different aggregation model hyperparameter sets. Specifically, each boxplot represents the performance range when fixing the feature extraction model by using 81 different hyperparameter sets for the aggregation model. We show this for three different datasets and for both AttentionMIL and TransMIL approaches. From the boxplots it can be seen that the performance fluctuation is large as well as the overlap between different feature extraction models. Furthermore, the boxplots show a similar trend when compared to using AUROC as metric.
Figure 6: The heatmap shows the performance of every aggregation model hyperparameter set for each feature extraction model according to the AP metric. The lighter the colour the higher the AP. The heatmaps are shown for all three datasets and for both AttentionMIL and TransMIL approaches. The red colored legend shows how the hyperparameter sets are ordered on the heatmap. For example, the lowest learning rate is used for the first third of the hyperparameter sets, followed by the next higher learning rate. The legend should facilitate detecting hyperparameters sets with consistently high performance. However, no clear patterns are visible in the heatmaps indicating that aggregation model hyperparameters need to be accounted for when comparing feature extraction models. We observe similar trends compared to when using AUROC as a metric.

### v3

@@ 2.4 Evaluation Datasets
Our study evaluates binary classification performance of feature extraction models across three distinct DP datasets. Thereby providing a more generalizable answer to our research questions. These datasets, comprising H&E-stained histopathology slides WSIs, allow us to assess each feature extractor under 162 different aggregation model configurations. This comprehensive approach, covering 7 feature extractors, 162 aggregation model configurations, and 3 datasets, culminates in a total comparison of $7\times 162\times 3=3402$ experimental configurations.
Performance metrics include the area under the receiver operating characteristic curve (AUROC) and average precision (AP), both ranging from 0 to 1. A higher AUROC indicates superior distinction between the positive and negative classes, while a higher AP reflects more accurate predictions of positive instances across all recall levels, effectively balancing precision and recall. Performance metrics are derived from the test set, using the aggregation model’s epoch with best validation score during training.
COO: Binary classification of cell of origin (COO). Each image contains the COO prediction label of activated B-cell like (ABC) or germinal center B-cell like (GCB) tumors in diffuse large B-cell lymphoma (DLBCL). 709 WSIs from two internal datasets were used. This data closely mirrors real-world data, since it is crucial to assess classification approaches in DP using such data and tasks. The WSIs (40$\times$ magnification) have been scanned by Ventana DP200 scanners. The artifact-free tissue tiles of this dataset were combined and randomly split into 70% training set, 15% validation set and 15% test set.
@@ 3 Results
This section addresses our initial inquiries, first assessing if a universally optimal aggregation model configuration exists for multiple feature extraction (foundation) models, and then comparing different feature extractors considering performance variability across aggregation model setups.
3.1 Aggregation Model Configuration Influence
Our analysis begins by evaluating the sensitivity of feature extractors, or foundation models, to various aggregation model configurations.
The heatmap in Figure 3 displays the classification performance across all aggregation model configurations. Trends are consistent across both AUROC and AP scores. The heatmap legend aids in identifying patterns, such as configurations with the lowest learning rate positioned on the left of each feature aggregator, which tend to yield lower performance in the COO dataset when using CTransPath but not for other features extractors. Analysis of these heatmaps reveals:
Figure 3: The heatmap shows the performance of every aggregation model configuration set for each feature extraction model. The red colored legend shows how the configurations are ordered on the heatmap.
Lack of a universal configuration: No single aggregation model configuration consistently outperforms across all feature extractors, as indicated by the absence of a uniformly bright column across models.
Dataset-specific configurations: Optimal configurations for a given feature extractor vary by dataset. While certain parameters like learning rate for AttentionMIL and the number of attention blocks for TransMIL show some dataset-specific importance, no definitive pattern emerges across datasets or models, suggesting the need for investigation of model-specific configurations.
These results highlight the need for evaluating a diverse range of configurations in the aggregation model. This approach would ascertain that any observed superiority of one feature extraction model over another is not simply attributed to the specific aggregation model setup selected.
@@ 3.2 Feature Extractor Comparison
Figure 4 diverges from the standard practice of showing a single outcome for a fixed aggregation model setup by presenting feature extractor model performance across all 162 configurations for various datasets. Through box plots, we observe substantial performance overlap among feature extraction models despite the variance across configurations. Key insights include:
Figure 4: Comparison of 7 feature extraction models across 162 different aggregation model configurations, which include 2 architectures with 81 parameters each.
Training on DP datasets is necessary: The DINOv2 model trained on natural images performs poorly compared to all other models, consistently for all the datasets.
Comparable Performance Across Model Sizes: The relatively small model $\text{BYOL}_{\text{TCGA}}$ matches the performance of larger ones, suggesting that larger models are not necessarily better for DP. This echoes Filiot et al.’s [8] findings that a ViT-B model can outperform a ViT-L model.
[…]
4 Conclusion
In this study, we challenge the prevailing methodology for comparing foundation models in digital pathology literature, demonstrating that it may yield misleading results. We show that due to the high sensitivity of feature extraction models to downstream aggregation model configurations, relying solely on a single aggregation model configuration can disproportionately favor certain feature extractor models while disadvantaging others. Hence, we propose evaluating foundation models across different configurations for fairer comparisons. Our comprehensive analysis, taking into account performance variations across multiple configurations of the aggregation model, reveals a considerable overlap in performance between different foundation or feature extractor models. Significantly, we find no universal aggregation model configuration that is uniformly effective for all feature extractors. Our work is limited though by only looking at classification tasks. In addition, the DINOv2 model we trained on digital pathology images might be subpar to other models due to computational and dataset limitations. Nevertheless, we believe this work will contribute to a more nuanced evaluation of foundation models that will help gain insight and further accelerate this rapidly evolving field.
Supplementary Material
Table S1: List of the feature extraction models used for evaluation with some of their key properties.
Learning Approach Model Details Model Size (Mio.) Embed. Size Dataset Content Dataset Size In-house training
[…]
Dino v2
Combination ViT-L/16 300 1024
TCGA + in-house
35 70 yes
Dino v2
Combination ViT-L/14 distilled 300 1024
LVD-142M
(natural images)
142 N/A no
iBOT Combination ViT-B 86 768 TCGA 43 6 no
CTransPath Contrastive Learning Swin Transformer 280 768 TCGA + PAIP 16 32 no
REMEDIS Contrastive Learning
BiT-M
(ResNet-152x2)
240 4096 TCGA 50 29 no
BYOL
Self-distillation ResNet50 26 2048 TCGA 2 30 yes
BYOL
Self-distillation ResNet50 26 2048
DLBCL, Herohe,
Camelyon16
2 1 yes
Figure S1: The heatmap shows the performance of every aggregation model configuration for each feature extraction model according to the AP metric. The lighter the color, the higher the AP. The heatmaps are shown for all three datasets. The red-colored legend shows how the aggregation model configurations are ordered on the heatmap. For example, the lowest learning rate is used for the first third of the configurations for each feature aggregator, followed by the next higher learning rate.
Table S2: Results for the different feature extractors across all configurations aggregated as mean $\pm$ standard deviation.
Feature extractor COO Herohe Camelyon
@@ 4 Conclusion
[…]
iBOT 0.76$\pm$ 0.05 0.79$\pm$ 0.05 0.72$\pm$ 0.06 0.61$\pm$ 0.07 0.88$\pm$ 0.13 0.88$\pm$ 0.14
CTransPath 0.63$\pm$ 0.06 0.71$\pm$ 0.05 0.70$\pm$ 0.06 0.59$\pm$ 0.07 0.76$\pm$ 0.12 0.75$\pm$ 0.14
REMEDIS 0.71$\pm$ 0.08 0.75$\pm$ 0.07 0.76$\pm$ 0.11 0.68$\pm$ 0.16 0.70$\pm$ 0.17 0.67$\pm$ 0.19
$\text{BYOL}_{\text{TCGA}}$ 0.71$\pm$ 0.04 0.77$\pm$ 0.04 0.77$\pm$ 0.05 0.68$\pm$ 0.08 0.84$\pm$ 0.12 0.84$\pm$ 0.13
$\text{BYOL}_{\text{CHC}}$ 0.77$\pm$ 0.07 0.82$\pm$ 0.06 0.79$\pm$ 0.05 0.69$\pm$ 0.07 0.87$\pm$ 0.09 0.87$\pm$ 0.09